# ABScribe: Rapid Exploration of Multiple Writing Variations in Human-AI Co-Writing Tasks using Large Language Models

### v1

@@ 3.1. Design Requirements
We surveyed literature on several key areas critical to our goal of facilitating the swift exploration of multiple writing variations using LLMs: the role and nuances of revision within the writing process (Sommers, 1980; Fitzgerald, 1987; MacArthur, 2018; Harris, 2017; Sharples, 2002; Flower and Hayes, 1981b); HCI design philosophies that emphasize the consideration of multiple ideas before evaluation (Tohidi et al., 2006; Greenberg and Buxton, 2008; Dow et al., 2010); principles on reification and reuse for designing visual interfaces (Beaudouin-Lafon and Mackay, 2000; Han et al., 2020b); and the latest research on utilizing LLMs in writing interfaces to foster experimentation and creativity (Yuan et al., 2022; Kim et al., 2023; Lee et al., 2022b). Based on this, we formulated an initial set of design requirements, which are summarized below.
@@
Requirement 1: Minimizing Task Workload while Exploring Multiple Variations of Text Drawing from HCI and traditional design principles, we emphasize the importance of exploring multiple ideas concurrently. Instead of refining a single solution to “get the design right”, these disciplines encourage the iteration and evaluation of multiple solutions in parallel to ultimately “get the right design” (Tohidi et al., 2006). We hypothesized that parallel exploration of text segments could aid writers during the revision process. We also considered that during the writing process writers frequently struggle with cognitive overload (McCutchen et al., 2009) and that even small demands on working memory can lead to decreased fluency (Ransdell et al., 2002). With this in mind, we hypothesized that an increase in writing variations could worsen this. As such, we aimed to design an editing interface that provides affordances for creating and comparing multiple writing variations without overwhelming the user.
@@
Requirement 2: Support visually-structured management of variations Documents can become quickly cluttered when trying to explore multiple variations of different text segments using current editing interfaces. Furthermore, the potential for LLMs to enhance writers’ ability to generate and revise multiple parallel variations of text further exacerbates the issue of clutter. Together, these factors highlight the need for a visually structured approach to manage variations. Our goal was to support writers in seamlessly integrating LLM-generated variations without cluttering the document or erasing existing content to retain the ability to simultaneously consider multiple variations. By presenting users with a range of variations, we give them the opportunity to select their favorites while simultaneously discarding less-favored alternatives (Greenberg and Buxton, 2008).
@@
Requirement 3: Support context-sensitive variation comparison and revision In a linear document editing interface, we found it difficult to maintain a sense of the surrounding text to situate new variations within existing context. This was particularly poignant when creating and comparing variations for smaller and embedded text segments - eg. A sentence mid-paragraph or paragraph mid-section - which disrupted the text flow. Maintaining text flow is crucial since writers need to engage with information processing tasks such as ensuring the document maintains cohesion which requires matching to surrounding text (McCulley, 1985) This becomes increasingly challenging as the document holds more and more variations of text segments. Our objective was to design an interface that allows writers to systematically evaluate these variations within context, eliminate less-favored options, and generate new iterations based on existing ones.
@@
Requirement 4: Supports revision-centric, reusable, and non-linear LLM usage Recognizing that revision is inherently nonlinear— with writers often revisiting earlier sections of a passage— and recursive, manifesting in repeated cycles throughout the writing process (Sommers, 1980; Fitzgerald, 1987), we aimed to align our LLM integrations with this fluid, iterative nature of revision. Our goal is for writers to be able to use LLMs to manipulate text segments of varying lengths and refine them as needed in a way that is natural to their non-linear and recursive process. To enable this, we draw inspiration from design principles for visual interfaces, focusing on reuse, polymorphism, and reification (Beaudouin-Lafon and Mackay, 2000), and regard LLM prompts as reusable, polymorphic commands that can be applied to targeted text segments of varying lengths, transforming them into first-class objects. Recent research on designing LLM-powered writing interfaces has highlighted the value of viewing components of the LLM generation pipeline as interactive objects in supporting iteration and experimentation (Kim et al., 2023). We adopt aspects of this approach in our design, such as reifying (Beaudouin-Lafon and Mackay, 2000) LLM prompts into reusable AI buttons, and turning text segments into interactive Variation Component objects or cells (Kim et al., 2023).
@@ 3.2. Interface Elements
We addressed these four design requirements by developing five interface elements using an iterative design process. To get the right design (Tohidi et al., 2006) the lead author iterated through multiple versions of each interface element, and tested them with a total of five pilot users during in-depth brainstorming and design-review sessions over six months.
@@
Variation Components: Users can select any part of the text and create interactive Variation Components that can hold multiple writing variations (Figure 2, Part 1). Newer variations can be added to an AB Component without overwriting existing variations (Figure 2 Parts 2 and 3), and without breaking the flow of the passage.
@@
Hover Buttons: Variations are represented using AB Hover Buttons that are dynamically placed above the active AB Component. Users can hover over each button to reveal the corresponding variation in the context of the surrounding passage (Figure 2, Parts 4 and 5). Moving the cursor away from the AB Hover Button reverts the AB Component back to the selected variation, allowing users to quickly compare between the selected and the hovered variation. Users can click on the Hover Button to select the variation and edit them in place, or discard it by clicking on the trash icon (Figure 2, Part 6).
@@
Variation Accordion: To help writers view multiple variations together, and navigate through them more easily, we pair the Hover Buttons with an accordion structure where each Variation Component has its own header, and all corresponding variations are stored underneath. Clicking on the variations in the accordion dynamically re-positions the Hover Buttons above the corresponding Variation Component, and vice versa, allowing users to manage multiple variations in a visually structured manner (Figure 3)
The Variation Components, Hover Buttons, and the Variation Accordion work together to address R1, R2 and R3. To tackle R4, we developed two interface elements.
@@
AI Buttons: Users can generate new variations by selecting an AB Component and typing instructions to the AI (Figure 3, Part 1). Instructions are automatically converted into labeled buttons (Figure 3, Part 2). The labels are generated using the LLM. As users experiment with newer variations using the AI, they create a set of custom AI Variation Buttons that they can reuse to apply to different parts of the passage, making these buttons reusable (Figure 3, Part 4). The prompts and labels for the buttons can be edited and improved over time. (Figure 4, Part 5). This allows writers to not only create a set of variations for a particular AB Component, but also design a set of buttons reflecting the kinds of variations they might want to generate for other parts of the text in the future, akin to a custom Swiss army knife for variations.
@@
AI Insert: Users can insert text from the AI model anywhere within the passage by writing instructions to the LLM in the following format: @ai ¡prompt¿. The text is generated and shown to the user in real-time, and they have the option to accept or discard the AI generated output, as well as revise the prompt to regenerate the output if it doesn’t match what the user is looking for (Figure 5).
@@ 4. Evaluating ABScribe
To validate our design, we conducted an within-subjects evaluation study where we compared ABScribe to a carefully constructed baseline interface. The Baseline interface featured rich-text editing capabilities commonly found in word processing software such as Google Docs and Microsoft Word, as well as a conversational AI assistant similar to ChatGPT, and the ability to incorporate AI generated text into the document without the need to copy and paste to represent the tighter AI integration available in modern AI editors such as Notion.AI444notion.ai. To minimize potential confounding factors due to tangential differences between the study conditions, we maintained the same overall layout for the common UI elements such as the width of the sidebar, placement and dimensions of buttons and text, font size, and color, and used the same underlying LLM model, GPT-4, to implement the generative AI features. We sought to answer the following research questions:
@@ 4.1. Participants
We recruited 12 writers (5 women, 7 men), aged 18 to 34, all of whom reported proficiency in reading and writing in English. They were screened for prior experience in a broad variety of both fiction and non-fiction writing genres. We included a range of prior experience levels of AI tool usage. This was done to ensure that our findings were not tied to specific genres or specific AI usage traits. Moreover, all participants had sufficient writing experience to be able to comment on the revision process and its applicability to different kinds of writing. See Table 1 for writer profiles.
@@ 4.2. Tasks
Each participant engaged in two guided writing tasks, which were randomly paired with the two counter-balanced study conditions. Our choice of tasks—writing an email and drafting a social media post—sought to provide an ecologically valid writing experience that fits timing constraints of the study and serves as a realistic use-case for LLMs. Such scenarios align with recent HCI studies on human-AI co-writing (Kim et al., 2023; Goodman et al., 2022). Additionally, commercial AI-assisted writing tools, such as Grammarly Go, Respondable, and Copy.ai, leverage AI for similar purposes.
Scenario Descriptions:
Then, using either the ABScribe or the Baseline interface, participants were asked to explore 8 variations (increasing or decreasing length, formality, word diversity, adding emojis, and two variations of their choosing) of 3 distinct text segments (title/subjectline, third sentence of the second paragraph, entire third paragraph), summing up to 24 variations. This approach ensured exploration of variations of consistent number, size and variety across study conditions, while affording some scope for creativity as, for two out of the eight variations for each segment, participants had the autonomy to craft variations based on their preferences.
@@ 4.3. Measures
To assess subjective task-workload, we used the widely used NASA-TLX (Hart, 2006) procedure with weighting. To quantify the specific aspects of the LLM-assisted revision process that we aimed to improve, we also asked participants to rate their agreement on a 7-point Likert scale (1 = Strongly Disagree, 7 = Strongly Agree), similar to prior work (Lee et al., 2022c; Kim et al., 2023; Lee et al., 2022b):
@@ 4.4. Procedure
Participants began by signing a consent form and completing a survey that captured demographic data, prior writing experience, and familiarity with AI-assisted writing tools. Conducted via video-conference, the entire study lasted approximately 1.5 hours. Participants accessed the prototypes through their web browsers, mirroring how they typically access popular AI editing tools like ChatGPT and Grammarly. Conducting the study online enabled us to engage a diverse group of writers beyond Canada. After a brief introduction outlining the study’s objective—to investigate various writing variations using LLM-integrated editing tools—participants undertook two 15-minute tasks. Before starting each task, we demonstrated how the tools functioned and gave participants an opportunity to try them, ensuring they felt comfortable. After each task, participants completed the NASA-TLX and 11 Likert-scale measures. These measures offered insights into the writers’ perceptions of the revision process and prompted them to reflect on specific aspects of the revision process that we seek to improve through our design. The evaluation concluded with a recorded 30-minute semi-structured user interview on their experience with each interface. As a token of our appreciation for their participation, each participant received 30 Canadian dollars.
@@ 4.5. Analysis
Our data comprised interview transcripts, task observation notes, and the NASA-TLX and Likert-scale ratings for each condition. We coded and analyzed the interview transcripts and task observation notes using reflexive thematic analysis (Braun and Clarke, 2019) through an inductive-deductive lens. The theory on revision-focused exploration of writing variations served as a pre-existing code guiding our interpretations.
In a within-subject design, we use pairwise one-sided t-tests to compare sum of scores of NASA-TLX and our Likert-scale measures on the revision process. T-test was shown to be robust for aggregated data of this kind. (de Winter and Dodou, 2010). Additionally, we aim to check for normality. We aimed to determine if ABScribe presented significant improvements over the baseline, prompting us to select a one-tailed test with hypothesis B ¡ A for task workload and B ¿ A for level of agreement on the efficacy of the revision process.
We performed an apriori power analysis for a pairwise one-sided t-test, showing that we can detect with power at least effect size with sample size participants for a significance level .

### v3

@@ 3.1. Design Requirements
We surveyed literature on several key areas critical to our goal of facilitating the swift exploration of multiple writing variations using LLMs: the role and nuances of revision within the writing process (Sommers, 1980; Fitzgerald, 1987; MacArthur, 2018; Harris, 2017; Sharples, 2002; Flower and Hayes, 1981b); HCI design philosophies that emphasize the consideration of multiple ideas before evaluation (Tohidi et al., 2006; Greenberg and Buxton, 2008; Dow et al., 2010); principles on reification and reuse for designing visual interfaces (Beaudouin-Lafon and Mackay, 2000; Han et al., 2020b); and the latest research on utilizing LLMs in writing interfaces to foster experimentation and creativity (Yuan et al., 2022; Kim et al., 2023a; Lee et al., 2022a; Buschek et al., 2021; Dang et al., 2023; Gero et al., 2022; Chung et al., 2022). Based on this, we formulated an initial set of design requirements, which are summarized below.
@@
Requirement 1: Minimize task workload when exploring multiple text variations. Drawing from HCI and traditional design principles, we emphasize the importance of exploring multiple ideas concurrently. Instead of refining a single solution to “get the design right”, these principles encourage the iteration and evaluation of multiple solutions in parallel to ultimately “get the right design” (Tohidi et al., 2006). We hypothesized that parallel exploration of text segments could aid writers during the revision process, in alignment with recent studies on the benefit of showing multiple parallel phrase suggestions on ideation (Buschek et al., 2021). We also considered that during the writing process writers frequently struggle with cognitive overload (McCutchen et al., 2009) and that even small demands on working memory can lead to decreased fluency (Ransdell et al., 2002). With this in mind, we hypothesized that an increase in writing variations could worsen this. As such, we aimed to design an editing interface that provides affordances for creating and comparing multiple writing variations without overwhelming the user.
@@
Requirement 2: Support visually-structured management of variations. Documents can quickly become cluttered when trying to work with multiple variations of different text segments using current editing interfaces. Furthermore, the potential for LLMs to enhance writers’ ability to generate and revise multiple parallel variations of text further exacerbates the issue of clutter. Together, these factors highlight the need for a visually structured approach to managing variations. Our goal was to support writers in seamlessly incorporating LLM-generated variations into the draft without cluttering the document or erasing existing content, and to retain the ability to simultaneously consider multiple variations. This is in alignment with prior studies that show that while users prefer being able to choose from multiple options (Dang et al., 2023), having all of these options can negatively impact writing efficiency (Buschek et al., 2021), indicating a need for option management support. Furthermore, having the ability to choose between options in an intuitive way can give users an opportunity to make more informed decisions as they compare and contrast between parallel ideas (Greenberg and Buxton, 2008; Tohidi et al., 2006; Buxton, 2010).
@@
Requirement 3: Support context-sensitive variation comparison and revision. In a typical, linear document editing interface, we found it difficult to maintain a sense of the surrounding text as we created and added variations. This was particularly noticeable when trying to create and compare variations for smaller and embedded text segments–e.g., a sentence mid-paragraph or paragraph mid-section–which disrupted the text flow. Maintaining text flow is crucial since writers need to engage with information processing tasks such as ensuring the document maintains cohesion, which requires matching segments to the surrounding text (McCulley, 1985). This becomes increasingly challenging as the document holds more and more variations of different text segments. Our objective was to design an interface that allows writers to systematically generate and edit multiple variations in context of the preceding and subsequent text in the document.
@@
Requirement 4: Supports revision-centric, reusable, and non-linear LLM usage. Recognizing that revision is inherently nonlinear–with writers often revisiting earlier sections of a passage–and recursive, manifesting in repeated cycles throughout the writing process (Sommers, 1980; Fitzgerald, 1987), we aimed to align our LLM integrations with this fluid, iterative nature of revision. Our goal was for writers to be able to use LLMs to manipulate text segments of varying lengths and refine them as needed in a way that is natural to their non-linear and recursive process. To enable this, we drew inspiration from design principles for visual interfaces, focusing on reuse, polymorphism, and reification (Beaudouin-Lafon and Mackay, 2000), and regarded LLM prompts as reusable, polymorphic commands that can be applied to targeted text segments of varying lengths. Recent research on designing LLM-powered writing interfaces has highlighted the value of viewing components of the LLM generation pipeline as interactive objects in supporting iteration and experimentation (Kim et al., 2023a). We adopt aspects of this approach in our design, such as reifying (Beaudouin-Lafon and Mackay, 2000) LLM prompts into reusable AI Modifiers, and turning text segments into interactive Variation Fields akin to cells (Kim et al., 2023a).
@@ 3.2. Interface Elements
We addressed these four design requirements by developing five interface elements using an iterative design process. To get the right design (Tohidi et al., 2006) the lead author iterated through multiple versions of each interface element, and tested them with five pilot users during in-depth brainstorming and design-review sessions over six months.
3.2.1. Design Process
The design process involved transitioning from low to high-fidelity prototyping via three overlapping processes: (i) Paper-Prototyping (Sefelin et al., 2003; Snyder, 2003): We began by considering an in-place editing interface as a foundational concept, and sketched various design elements that could be integrated into this interface to enhance exploration of variations; (ii) Cognitive Walkthroughs (Lewis and Wharton, 1997; Mahatody et al., 2010): We used informal cognitive walkthroughs to rapidly iterate on our sketches and identify which potential solutions were intuitive; (iii) Web-based Prototyping (Walker et al., 2002): We then developed functional versions of our intuitive design solutions in a high-fidelity prototype. There was considerable iteration between these stages. For example, if we saw issues in the high-fidelity, we would return to paper prototyping to devise refinements. Through this process, we developed five interface elements:
3.2.2. Variation Fields
@@
Users can select any part of the text and create interactive Variation Fields that can hold multiple writing variations (Figure 2, Part 1). Newer variations can be added to an Variation Field without overwriting existing variations (Figure 2 Parts 2 and 3) and without breaking the flow of the passage.
3.2.3. Popup Toolbar
@@
Variations are represented using buttons residing within a Popup Toolbar placed above the active Variation Field. The Popup Toolbar dynamically moves with the text caret, and users can peek at different variations by hovering over the corresponding buttons to view them in the context of the surrounding passage (Figure 2, Parts 4 and 5). Moving the cursor away from a toolbar button reverts the Variation Field back to the selected variation, allowing users to quickly compare between the selected and the hovered variations. Users can click on a button to select the variation and edit them in place, or discard it by clicking on the trash icon (Figure 2, Part 6).
3.2.4. Variation Sidebar
@@
To help writers view multiple text variations together, and navigate through them more easily, we pair the Popup Toolbar with a sidebar organized in an accordion structure where each Variation Field has its own header, and all corresponding variations are stored underneath. Clicking on the variations in the sidebar dynamically re-positions the Popup Toolbar above the active Variation Field, and vice versa, allowing users to navigate through multiple variations in a visually structured manner (Figure 3). Users can show or hide the sidebar from the editor menu.
The Variation Fields, Popup Toolbar, and the Variation Sidebar work together to address R1, R2 and R3. Note that the Popup Toolbar and Variation Fields have no AI capability, but support text revision by the user. To tackle R4, we developed two interface elements.
3.2.5. AI Modifiers
@@
Users can generate new variations by selecting a Variation Field and typing instructions to the AI (Figure 4, Part 1 and 2). Instructions are automatically converted into labeled buttons (Figure 4, Part 4). The labels are generated using the LLM. As users experiment with newer variations using the AI, they create a set of custom AI Modifiers that they can apply to different parts of the passage, making them reusable (Figure 4, Part 5 and 6). The prompts and labels for the buttons can be edited and improved over time (Figure 4, Part 7 and 8). This allows writers to not only create variations for a particular Variation Field, but also design buttons reflecting the kinds of variations they might want to generate for other parts of the text in the future, akin to a custom Swiss army knife for variations. As users add more variations, the interface organizes them into a scrollable list with a maximum height, so users do not run out of room. A more complete implementation could explore more sophisticated ways to group and search through AI Modifiers.
3.2.6. AI Drafter
@@
Users can insert text from the AI model anywhere within the passage by prompting the system to invoke LLM usage in the following format: @ai ¡prompt¿, and pressing enter. The text is generated and shown to the user in real-time, and they have the option to accept or discard the AI generated output, as well as revise the prompt to regenerate the output if it does not match what the user is looking for (Figure 5).
@@ 4. Evaluating ABScribe
To validate our design, we conducted a within-subjects evaluation study where we compared ABScribe to a carefully constructed baseline interface. Refer to Appendix A to see a screenshot. The Baseline interface featured the same rich-text editing capabilities as ABScribe without the ability to create and modify Variation Fields using the Popup Toolbar, Variation Sidebar and AI Modifiers. Instead, it featured a conversational AI assistant similar to ChatGPT, and the ability to insert AI generated text directly into the document as some modern AI editors such as Notion AI555notion.ai have that capability. To minimize potential confounding factors arising from tangential differences between the study conditions, we maintained the same overall layout for common UI elements. This included the width of the sidebar, placement and dimensions of buttons and text fields, font size, and color. Additionally, we used the same underlying LLM model, GPT-4, to implement the generative AI features for both conditions.
@@ 4.1. Participants
We recruited 12 writers (5 women, 7 men, aged 18 to 34) through flyers and study invitation messages on social networking services, such as LinkedIn and Facebook. All participants reported proficiency in reading and writing in English and were screened for prior experience in a broad variety of both fiction and non-fiction writing genres. We included participants with a range of prior experience levels in AI tool usage. This ensured that our findings were not limited to specific genres or AI usage patterns. All participants had ample writing experience, enabling them to effectively comment on the revision process and its relevance across various writing forms. Refer to Table 1 for writer profiles. Participants were compensated with 30 CAD. The institution’s research ethics board has approved the protocol.
@@ 4.2. Tasks
Each participant engaged in two guided writing tasks which were randomly paired with the two counter-balanced study conditions. Our choice of tasks—writing an email and a social media post—aimed to provide an ecologically valid writing experience that fits the timing constraints of our study and served as realistic use-cases for LLMs, in alignment with recent HCI studies on human-AI co-writing (Kim et al., 2023a; Goodman et al., 2022; Buschek et al., 2021), and commonly advertised applications of commercial AI-assisted writing tools, such as Grammarly Go, Respondable, and Copy.ai. We selected scenarios by considering situations that: (i) would be easy for users to imagine, such as emailing a professor or seeking a job as a copywriter; (ii) offer opportunities for exploring variations, such as devising multiple alternatives for a subject line or altering a sentence to maximize reader engagement. We wrote and tested the prompts to ensure the output was reasonable and of consistent length. For example, after initial rounds of testing, we found that the output generated by the AI was too long, and so we added ‘Keep it within three paragraphs’ to both prompts. We asked participants to generate a draft using the same prompts for all scenarios. This approach was designed to keep their focus on exploring and organizing variations, rather than engaging in prompt engineering on the initial draft. It also helped maintain relative consistency in the generated output across participants and study conditions. Refer to Appendix B for additional details on the task scenario descriptions and prompts.
Then, with either ABScribe or the Baseline interface, participants were asked to use AI to explore eight variations (increasing or decreasing length, formality, word diversity, adding emojis, and two variations of their choosing) of three distinct text segments (title/subjectline, third sentence of the second paragraph, entire third paragraph), summing up to twenty-four variations. This approach ensured variation exploration of consistent number, size and variety across study conditions, while affording some scope for creativity as, for two out of the eight variations for each segment, participants had the autonomy to craft variations based on their preferences. We selected the variations by considering potential modifications writers might want to apply to their drafts, including changes in length, tone, and word choice. We then tested these modifications to ensure their feasibility with both ABScribe and the baseline interface.
@@ 4.3. Measures
To assess subjective task-workload, we used the widely used NASA-TLX (Hart, 2006) procedure with weighting. To quantify the specific aspects of the LLM-assisted revision process that we aimed to improve, we also asked participants to rate their agreement on a 7-point Likert scale (1 = Strongly Disagree, 7 = Strongly Agree), similar to prior work (Lee et al., 2022b; Kim et al., 2023a; Lee et al., 2022a):
@@ 4.4. Procedure
Participants began by signing a consent form and completing a survey that captured demographic data, prior writing experience, and familiarity with AI-assisted writing tools. The entire study lasted approximately 1.5 hours, and was conducted via video-conferencing software. Participants accessed the prototypes through their web browsers, mirroring how they typically access popular AI editing tools like ChatGPT and Grammarly. Conducting the study online enabled us to engage a diverse group of writers beyond Canada. After a brief introduction outlining the study’s objective—to investigate the exploration and organization of multiple writing variations using LLM-integrated editing tools—participants undertook the two 15-minute writing tasks. Before starting each task, we demonstrated how the tools functioned and gave participants an opportunity to try them, ensuring they felt comfortable. After each task, participants completed the NASA-TLX assessment and 11 Likert-scale measures. These measures offered insights into the writers’ perceptions of the revision process and prompted them to reflect on specific aspects of the process that we sought to improve through our design. The evaluation concluded with a recorded 30-minute semi-structured user interview on their experience with each interface.
@@ 4.5. Analysis
Our data comprised interview transcripts, task observation notes, and the NASA-TLX and Likert-scale ratings for each condition. We coded and analyzed the interview transcripts and task observation notes using reflexive thematic analysis (Braun and Clarke, 2019) through an inductive-deductive lens. The theory on revision-focused exploration of writing variations served as a pre-existing code guiding our interpretations. In a within-subjects design, we use pairwise one-sided t-tests to compare the sum of scores of NASA-TLX and our Likert-scale measures on the revision process. T-test was shown to be robust for aggregated data of this kind (de Winter and Dodou, 2010). We aimed to determine if ABScribe presented significant improvements over the baseline interface, prompting us to select a one-tailed test with hypothesis for task workload and for level of agreement on the efficacy of the revision process. We performed an a priori power analysis for a pairwise one-sided t-test, showing that we can detect an effect size of at least with power and a sample size of participants for a significance level of .